# DPCore: Dynamic Prompt Coreset for Continual Test-Time Adaptation

### v3

@@ 5 Conclusion
Impact Statements
@@
Appendix A Related Works
Test-Time Adaptation (TTA) (Wang et al., 2021; Iwasawa & Matsuo, 2021; Liang et al., 2020; Gao et al., 2023; Schneider et al., 2020; Mirza et al., 2022, 2023; Niu et al., 2024; Liang et al., 2023; Sun et al., 2020) enhances pre-trained model performance using unlabeled data at test time, without access to the original training phase. TTA techniques fall into two main categories based on their use of source data. The first adjusts models through self-supervised losses like entropy minimization (Wang et al., 2021; Niu et al., 2023) and consistency maximization (Wang et al., 2022a; Liu et al., 2023b, 2024). The second involves preliminary steps using source data: either by extracting source characteristics such as statistics or features (Mirza et al., 2023; Niu et al., 2024; Zhang et al., 2024)) or by warming up injected parameters on source data before adaptation (Lee et al., 2024; Song et al., 2023; Gan et al., 2023; Liu et al., 2023b; Gao et al., 2023).
Crucially, both approaches are source-free adaptation since they operate without source data during adaptation. While our method falls into the second category, it significantly reduces source dependency compared to existing methods: DPCore requires only 300 unlabeled source examples to compute statistics, in contrast to DePT (Gao et al., 2023), ViDA (Liu et al., 2023b), VDP (Gan et al., 2023) and BeCoTTA (Lee et al., 2024) which need the entire source dataset for warm-up. Moreover, as demonstrated in Fig. 5c, our approach maintains effectiveness even when source data is completely unavailable.
Continual Test-Time Adaptation (CTTA) (Wang et al., 2022a; Lee et al., 2024; Niu et al., 2022; Niloy et al., 2024; Liu et al., 2023a; Boudiaf et al., 2022; Gong et al., 2022; Yuan et al., 2023) tackles the challenge of non-stationary target domains. This paradigm was first introduced in (Wang et al., 2022a), referred to as CSC in this work and shown in the top of Fig. 1, where domains change in a structured manner with clear boundaries and uniform lengths. While several works (Niu et al., 2023; Boudiaf et al., 2022; Gong et al., 2022) explore TTA challenges such as mixed domains and label imbalance, they primarily focus on the static domain. (Yuan et al., 2023) extends this exploration to the CSC setting, investigating label imbalance within batches from changing domains. More recently, (Lee et al., 2024) introduced scenarios with continual gradual shifts, where domain boundaries become blurred and transitional batches may contain mixed-domain data, though the overall domain progression still follows the CSC pattern.
Existing methods often overlook frequent domain changes and irregular durations, which are critical in real-world applications. To address this gap, we propose the Continual Dynamic Change (CDC) setting, where domains change in a more realistic, dynamic manner. Traditional CTTA methods struggle in this setting, as demonstrated in Fig. 4b. Our DPCore is specifically designed to mitigate the effect of diverse domain change patterns and frequencies.
Appendix B Baselines
@@
Appendix C Additional Details of DPCore
C.1 Visual Prompt Adaptation in TTA
In the context of TTA where labels are unavailable, prompts can be optimized through consistency maximization, entropy minimization, or unsupervised distribution alignment (Niu et al., 2024; Sun et al., 2023; Gao et al., 2023; Zhang et al., 2024). We employ distribution alignment due to its proven effectiveness and computational simplicity (Ben-David et al., 2006; Mehra et al., 2024).
@@ C.1 Visual Prompt Adaptation in TTA
$\bm{p}^{*}=\mathop{\arg\min}\limits_{\bm{p}}\;d(\mathcal{Z}^{S},\mathcal{Z}^{T}_{t}(\bm{p}))$ (8)
@@ C.3 Implementation Details
C.4 Comprehensive Analysis of DPCore
@@
From Proposition C.1, correct assignment at time $t$ depends only on well-separatedness and correct assignments at $t-1$, not on batch order. For the second claim, consider batches $B^{(1)},B^{(2)},...$ are the batches assigned to cluster $j$ in the order they are processed. With $\alpha=\frac{1}{|G^{j}|}$, updates yield: $\bm{\mu}^{j}=\bm{\mu}(B^{(1)})$, $\bm{\mu}^{j}=\frac{1}{2}(\bm{\mu}(B^{(1)})+\bm{\mu}(B^{(2)}))$, $\bm{\mu}^{j}=\frac{1}{3}(\bm{\mu}(B^{(1)})+\bm{\mu}(B^{(2)})+\bm{\mu}(B^{(3)}))$, and so on. This running average is independent of batch order. ∎
Proposition C.3.
@@
From Proposition C.2, cluster assignment is independent of batch order. For batch $B_{t}$, two cases arise: 1) If $B_{t}$ belongs to an existing cluster $i$, $\bm{p}_{t}$ is learned starting from existing core prompt $\bm{p}^{i}$. Since $\bm{p}^{i}$ is order-independent, $\bm{p}_{t}$ is also order-independent. 2) If $B_{t}$ creates a new cluster, $\bm{p}_{t}$ is learned from scratch, making it inherently order-independent. ∎
@@
Appendix D Additional Results
D.1 Performance of Single Domain-learned Prompt
@@
[…]
D.2 Results for 10 Random Orders in CSC
@@
D.3 Results for CIFAR10/100-to-CIFAR10/100C
@@
Similarly, for CIFAR100-to-CIFAR100C task (results in Tables 12 and 13), DPCore demonstrates strong performance across both settings. In CSC, DPCore achieves a 10.3% improvement over the source model and outperforms ViDA by 2.2%, with notable improvements across all corruption types, particularly on Glass (44.1% vs 60.5% source error rate) and Contrast corruptions (13.2% vs 34.8% source error rate). The CDC setting presents a greater challenge, where several methods (CoTTA, VDP) perform worse than the source model, and ViDA’s improvement drops from +8.1% to +5.6%. In contrast, DPCore maintains robust performance with an 8.0% improvement over the source model, demonstrating its effectiveness.
@@
[…]
D.4 Comprehensive Results for Cityscapes-to-ACDC
@@
Appendix E Discussion on CTTA Setting.
@@
[…]
E.1 CDC with Different Dirichlet Distributions
@@
E.2 CDC with Other Distributions
@@
[…]
E.3 Comparison with Other CTTA Settings
@@
Second, the gradual domain change setting proposed in (Lee et al., 2024) assumes blurred domain boundaries, where batches near transitions contain mixed-domain data. While this realistic assumption differs from traditional CTTA methods, it still follows the CSC pattern where mixed batches constitute only a small portion of all data due to long domain durations. Although we haven’t directly evaluated DPCore in this setting, our proposed variant DPCore-B (detailed in Section F.1) demonstrates effective handling of mixed-domain batches, suggesting potential applicability to gradual domain changes.
@@
Appendix F Additional Ablation Studies
@@
F.1 DPCore with Small Batch Size
@@
F.2 DPCore without Access to Source Data
In this section, we explore an extreme scenario where source data is completely inaccessible, further demonstrating DPCore’s practicality in real-world applications. Following the insights from (Mehra et al., 2024), we propose using public datasets that share similar or identical label spaces. For instance, in ImageNet-to-ImageNet-C or CIFAR10-to-CIFAR10C tasks, we leverage STL10 (Coates et al., 2011), which shares label space with CIFAR10 despite being distinct from ImageNet.
@@ F.2 DPCore without Access to Source Data
F.3 Details in Computational Complexity
@@
Appendix G Limitations

### v4

@@ 6 Conclusion
Acknowledgment
We thank the anonymous reviewers for their insightful comments and suggestions. We are also grateful to Evan Shelhamer for his thoughtful feedback on connecting our work to the broader continual test-time adaptation literature. This work was supported by the NSF EPSCoR-Louisiana Materials Design Alliance (LAMDA) program #OIA-1946231.
Impact Statement
@@ Acknowledgment
Appendix A Related Works and Baselines
A.1 Continual Learning (CL)
CL (Rebuffi et al., 2017; De Lange et al., 2021; Farajtabar et al., 2020; Rolnick et al., 2019; Li & Hoiem, 2017; Kirkpatrick et al., 2017; Zenke et al., 2017; Silver & Mercer, 2002; Bobu et al., 2018; Mai et al., 2022, 2021; Shim et al., 2021) addresses catastrophic forgetting by enabling models to retain learned information while acquiring new knowledge. This challenge is relevant to to TTA, especially in Continual Dynamic Change (CDC) environments, where adapting to new target domains often degrades performance on previously seen domains. To address these limitations, our approach integrates principles from CL into the TTA framework. We employ a dynamically updating coreset that preserves knowledge across visited domains, alongside a weighted updating mechanism to reduce error accumulation.
A.2 Baselines
@@
RDumb [^4] https://github.com/oripress/CCC (Press et al., 2023) employs periodic resets to the original pretrained weights at regular intervals to prevent the accumulation of adaptation errors that lead to performance collapse in continual test-time adaptation.
@@
Appendix B Additional Details of DPCore
B.1 Visual Prompt Adaptation in TTA
In the context of TTA where labels are unavailable, prompts can be optimized through consistency maximization, entropy minimization, or unsupervised distribution alignment (Niu et al., 2024; Sun et al., 2023; Gao et al., 2023; Zhang et al., 2024). We employ distribution alignment due to its proven effectiveness and computational simplicity (Ben-David et al., 2006; Mehra et al., 2024a).
@@ B.1 Visual Prompt Adaptation in TTA
$\bm{p}^{*}=\mathop{\arg\min}\limits_{\bm{p}}\;d(\mathcal{Z}^{S},\mathcal{Z}^{T}_{t}(\bm{p})),,$ (8)
@@ B.3 Implementation Details
B.4 Comprehensive Analysis of DPCore
@@
From Proposition B.1, correct assignment at time $t$ depends only on well-separatedness and correct assignments at $t-1$, not on batch order. For the second claim, consider batches $B^{(1)},B^{(2)},...$ are the batches assigned to cluster $j$ in the order they are processed. With $\alpha=\frac{1}{|G^{j}|}$, updates yield: $\bm{\mu}^{j}=\bm{\mu}(B^{(1)})$, $\bm{\mu}^{j}=\frac{1}{2}(\bm{\mu}(B^{(1)})+\bm{\mu}(B^{(2)}))$, $\bm{\mu}^{j}=\frac{1}{3}(\bm{\mu}(B^{(1)})+\bm{\mu}(B^{(2)})+\bm{\mu}(B^{(3)}))$, and so on. This running average is independent of batch order. ∎
Proposition B.3.
@@
From Proposition B.2, cluster assignment is independent of batch order. For batch $B_{t}$, two cases arise: 1) If $B_{t}$ belongs to an existing cluster $i$, $\bm{p}_{t}$ is learned starting from existing core prompt $\bm{p}^{i}$. Since $\bm{p}^{i}$ is order-independent, $\bm{p}_{t}$ is also order-independent. 2) If $B_{t}$ creates a new cluster, $\bm{p}_{t}$ is learned from scratch, making it inherently order-independent. ∎
@@
Appendix C Additional Results
C.1 Performance of Single Domain-learned Prompt
@@
DPCore (Proposed) 42.2 38.7 39.3 47.2 51.4 47.7 46.9 39.3 36.9 37.4 22.0 44.4 45.1 30.9 29.6 39.9 +15.9
C.2 Results for 10 Random Orders in CSC
@@
C.3 Results for CIFAR10/100-to-CIFAR10/100C
@@
Similarly, for CIFAR100-to-CIFAR100C task (results in Tables 13 and 14), DPCore demonstrates strong performance across both settings. In CSC, DPCore achieves a 10.3% improvement over the source model and outperforms ViDA by 2.2%, with notable improvements across all corruption types, particularly on Glass (44.1% vs 60.5% source error rate) and Contrast corruptions (13.2% vs 34.8% source error rate). The CDC setting presents a greater challenge, where several methods (CoTTA, VDP) perform worse than the source model, and ViDA’s improvement drops from +8.1% to +5.6%. In contrast, DPCore maintains robust performance with an 8.0% improvement over the source model, demonstrating its effectiveness.
@@
DPCore(Proposed) 24.1 20.6 24.5 13.9 26.5 14.2 13.2 13.4 10.2 12.8 5.0 22.5 16.8 20.1 18.5 17.1 +11.1
C.4 Comprehensive Results for Cityscapes-to-ACDC
@@
Appendix D Discussion on CTTA Setting.
@@
DPCore (Proposed) 42.7 40.4 42.2 57.4 61.0 51.1 52.4 35.8 41.0 38.8 22.1 55.4 48.3 31.1 28.1 43.2 +12.6
D.1 CDC with Different Dirichlet Distributions
@@
D.2 CDC with Other Distributions
@@
DPCore 39.9 (+15.9) 43.9 (+12.0) 42.1 (13.7) 42.0 (+13.9)
D.3 Comparison with Other CTTA Settings
@@
Second, the gradual domain change setting proposed in (Lee et al., 2024) assumes blurred domain boundaries, where batches near transitions contain mixed-domain data. While this realistic assumption differs from traditional CTTA methods, it still follows the CSC pattern where mixed batches constitute only a small portion of all data due to long domain durations. Although we haven’t directly evaluated DPCore in this setting, our proposed variant DPCore-B (detailed in Section E.1) demonstrates effective handling of mixed-domain batches, suggesting potential applicability to gradual domain changes.
@@
Appendix E Additional Ablation Studies
@@
E.1 DPCore with Small Batch Size
@@
E.2 DPCore without Access to Source Data
In this section, we explore an extreme scenario where source data is completely inaccessible, further demonstrating DPCore’s practicality in real-world applications. Following the insights from (Mehra et al., 2024a), we propose using public datasets that share similar or identical label spaces. For instance, in ImageNet-to-ImageNet-C or CIFAR10-to-CIFAR10C tasks, we leverage STL10 (Coates et al., 2011), which shares label space with CIFAR10 despite being distinct from ImageNet.
@@ E.2 DPCore without Access to Source Data
E.3 Details in Computational Complexity
@@
Appendix F Discussions and Limitations
F.1 Potential Failure Cases
Our method, DPCore, is designed to be robust against scenarios such as Overlapping Distributions, Noisy or Small Domain Shifts, and Inconsistent Weighting Effects being inherent failure cases. The core objective of DPCore is not to identify an identical number of prompts as domains or to pinpoint every distinct domain. Instead, it aims to update the same prompt for groups of similar domains. This is supported by our findings (Sec. 4.3 and Appendix C.1) that a prompt learned for one domain (e.g., Gaussian Noise) can be effective for a similar domain (e.g., Shot Noise) and can be further improved with minor updates. Due to the use of distance-induced weights, a new domain can be effectively represented as a decomposition of existing domains, allowing the weighted prompt to perform well even if derived from different domains, as domain similarity is assessed by distance and prompts are learned through distance minimization. A more significant potential failure case arises when each test batch contains data from multiple domains. In this situation, batch statistics would become unstable, potentially leading DPCore to treat each batch as a new, distinct domain. This would significantly reduce efficiency, as the algorithm would need to learn prompts from scratch for each batch, a scenario analogous to our experimental setup in Table 3 Exp-2 (where prompts were learned from scratch for each batch, though each batch was from a single domain). Such multi-domain batches pose a challenge to most CTTA methods, which typically assume single-domain test batches. We identify this as an avenue for future investigation.
F.2 Limitations